OpenStudy (anonymous):

[0,pi] cot(2x)=-1/sqrt(3) solve for x

OpenStudy (lgbasallote):

\[\cot (2x) = -\frac 1{\sqrt 3}\] take arc cot of both sides \[2x = \cot^{-1} \left(-\frac 1{\sqrt 3} \right)\] divide both sides by 2 \[\Large x = \frac{\cot^{-1} \left(-\frac 1{\sqrt 3} \right)}2\] simplify to get the final answer does that help?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well, I have to solve for x on that interval.

OpenStudy (lgbasallote):

yes. simplify what i just wrote and you'll get x...

OpenStudy (lgbasallote):

tehn just pick the values within [0,pi]